### v1

# Impact chronology of leftover planetesimals

Abstract
Context. After the formation of the Moon the terrestrial planets were pummelled by impacts from planetesimals left over from terrestrial planet formation. Most lunar crater chronologies are fit with an exponentially decreasing impact rate with an e-folding time of about 150 Myr. Dynamical simulations consisting of leftover planetesimals and the asteroid belt should reproduce this impact rate.
Aims. This work attempts to reproduce the impact rates set by modern crater chronologies using leftover planetesimals from three different dynamical models of terrestrial planet formation.
Methods. I ran dynamical simulations for 1 billion years using leftover planetesimals from the Grand Tack, Depleted Disc and Implantation models of terrestrial planet formation with the CPU version of the Gravitational ENcounters with GPU Acceleration (GENGA) N-body integrator. I fit the cumulative impacts on the Earth and Mars as well as the fraction of remaining planetesimals using a function that is a sum of exponentials with different weighing factors and e-folding times.
Results. Most fits require three or four terms. The fitted timescales cluster around million years (Myr), Myr, Myr and Myr. I attribute them to dynamical losses of planetesimals through different mechanisms: high-eccentricity Earth crossers and the secular resonance, Earth crossers, Mars crossers, and objects leaking on to Mars crossing orbits from beyond Mars. I place a constraint on the initial population using the known Archean terrestrial spherule beds, and I conclude that the Archean impacts were mostly created by leftover planetesimals. The inferred mass in leftover planetesimals at the time of the Moon’s formation was about 0.015 Earth masses.
Conclusions. The third time constant, , is comparable to that of modern crater chronologies. As such, the crater chronologies are indicative of impacts by an ancient population of Mars crossers. The initial perihelion distribution of the leftovers is a major factor in setting the rate of decline: to reproduce the current crater chronologies the number of Earth crossers at the time of the Moon’s formation had to be at most half of the Mars crossers. These results together place constraints on dynamical models of terrestrial planet formation.
Key Words.:
Planets and satellites: dynamical evolution and stability;1 Introduction
In traditional dynamical models the terrestrial planets grew from a coagulation of planetesimals into protoplanets, which continued to remain submerged in a swarm of planetesimals for millions of years. This system eventually evolved into a giant impact phase, wherein the protoplanets collided with the planetesimals and with each other, which ultimately lead to the terrestrial planets. There exist many different dynamical simulations to explain this history (e.g. Chambers, 2001; Raymond et al., 2006, 2009; O’Brien et al., 2006; Walsh & Levison, 2019; Woo et al., 2021); a review of this can be found in Morbidelli et al. (2012), and a brief discussion of each model’s outcomes are given in Lammer et al. (2021).
One key feature that all of the models share is that the terrestrial planets experience a protracted history of late accretion subsequent to planet formation. After core formation and initial separation of the silicate reservoirs (crust/mantle), left-over planetesimals on planet-crossing orbits were consumed by the terrestrial bodies as substantial mass supplements (Wetherill, 1977; Day et al., 2012). This took the form of impact bombardment by comets (Gomes et al., 2005), planetesimals remaining from primary accretion (Bottke et al., 2007), and asteroids from the main belt (Minton & Malhotra, 2010; Nesvorný et al., 2017) and the purported E-belt (Bottke et al., 2012). This late accretion thermally, structurally and chemically modified solid surfaces, most of which is directly visible as craters today.
In the current Solar System, small objects that venture near the Earth – so called Near-Earth Objects, or NEOs – consist of asteroids and comets; the latter are mostly the Jupiter-family Comets (JFCs). The asteroids that venture near the Earth – the Near-Earth Asteroids, or NEAs – are typically injected onto Earth-crossing orbits from the main asteroid belt through the secular resonance near 2 au, the 3:1 Jovian mean-motion resonance near 2.5 au, and the Mars-crossing population (e.g. Gladman et al., 1997; Morbidelli & Gladman, 1998; Bottke et al., 2002; Granvik et al., 2018). Main belt asteroids are driven towards Mars-crossing orbits by numerous weak mean-motion, secular and three-body resonances (Migliorini et al., 1998; Nesvorný & Morbidelli, 1998; Morbidelli & Nesvorný, 1999), as well as the Yarkovsky effect (Farinella et al., 1998; Farinella & Vokrouhlický, 1999; Vokrouhlický & Farinella, 1998, 1999, 2000). From there they are rapidly driven to Earth-crossing orbits, and their population declines with a half life ranging from million years (Myr) for objects in the and 3:1 resonances (Gladman et al., 1997), to 60 Myr for the low-inclination Mars-crossers, to over 100 Myr for the high-inclination Mars-crossers (Migliorini et al., 1998; Michel et al., 2000; Bottke et al., 2002; Fernández & Helal, 2023).
It is not obvious, however, that in the early Solar System these same timescales modulated the impact rate on the terrestrial planets. During this early epoch the impact rate was dominated by planetesimals left over from terrestrial planet formation, whose orbital distribution was very different from that of the main belt today (e.g Morbidelli et al., 2018; Brasser et al., 2020; Nesvorný et al., 2022). A constraint on the impact rate comes from crater chronology on the Moon and Mars, which for the first one billion years (Gyr) is fit well by a simple exponential decline with an e-folding time of 145-150 Myr for the Moon (Neukum et al., 2001; Werner et al., 2023), and 200 Myr for Mars (Werner et al., 2014; Werner, 2019). Even though the e-folding time of the Neukum and Werner chronologies are similar, the absolute crater density of the Werner chronology is much lower, resulting in the Werner chronology predicting ages up to 200 Myr older, and the latter requiring a much less massive impacting population. Dynamical simulations of leftover planetesimals, however, yield decline rates that vary greatly and are not necessarily consistent with that of the published crater chronologies (Morbidelli et al., 2018; Brasser et al., 2020; Nesvorný et al., 2022).
The reason for these discrepancies could be that the initial conditions of the planetesimals at the time of the Moon-forming event were not consistent with the population that produced the current craters on the Moon and Mars. Both Morbidelli et al. (2018) and Brasser et al. (2020) relied on leftover planetesimals from terrestrial planet simulations in the framework of the Grand Tack model (Walsh et al., 2011), while Nesvorný et al. (2022) relied on results from the annulus model (Hansen, 2009). Yet none of these initial conditions appear very satisfactory in reproducing cratering rate advocated by the established crater chronologies and many different models of terrestrial planet formation are present in the literature. These models likely produce different leftover populations.
In this work I investigate what the impact rate is of leftover planetesimals of three models of terrestrial planet formation simulations in my database, and whether these are or can be made consistent with modern crater chronologies. I also place constraints on what the orbital distribution of this population might have been, which of these models best reproduces this distribution, and what these constraints imply for models of terrestrial planet formation.
2 Numerical simulations: initial conditions and methods
This work relies on long-term simulations of massless test particles left over from simulations of terrestrial planet formation. Following Brasser et al. (2020), the initial conditions for this work were taken from a database of terrestrial planet formation simulations in my possession based on three models: the Grand Tack (GT) model (Walsh et al., 2011; Brasser et al., 2016), the Implantation (ABI) model (Brasser, 2025) and the Depleted Disc (DD) model (Izidoro et al., 2014; Mah & Brasser, 2021). I took a snapshot of the position and velocity vectors of the planetesimal population 60 Myr after the start of the simulations, which approximately coincides with 4.5 Ga. From these data I filtered out all planetesimals that had a perihelion distance au (2.1 au for DD), and an aphelion distance au so that they do not immediately venture near Jupiter. The initial values of semi-major axis (), eccentricity () and inclination () are shown in the top three panels of Fig. 1. The grey, orange, blue and red lines denote the required eccentricity to cross the orbits of Mercury, Venus, Earth and Mars, respectively. The colour coding is a proxy for the inclination, indicated by the colourbar. The bottom three panels show the cumulative distributions of perihelion, inclination and semi-major axis for all three models. All cumulative distributions are similar to either logistic or (offset) Rayleigh distributions. Specifically, the cumulative perihelion and inclination distributions are (offset) Rayleigh, and the semi-major axis distributions are logistic. These distributions are given by
where e is the base of the natural logarithm. The fitting constants for the perihelion are au, with the widest for DD, and with equal to the minimum of about 0.35 au for ABI, 0.41 and DD and 0.1 au for GT. For the inclination I obtain with . For the semi-major axis au for GT and 1.65 au for DD and ABI, and au for ABI and DD, and 0.32 au for GT. Uncertainties in the fit are for the inclination and au for and .
The GT model initially has the most Earth crossers (70% of all planetesimals), while the DD model has the least (only 30%). The ABI model has the highest average inclination, while this is lowest for the DD model. The planetesimals in the GT model are generally closer to the Sun than in the ABI and DD models, for which the cumulative semi-major axis distributions are very similar up to 2 au. Given these different initial conditions the expectation is that the dynamical evolution of all three populations will evolve somewhat differently.
The initial number of planetesimals from each set of terrestrial planet formation simulations was between 9 000 (9k) and 35 000 (35k), so I created clones of the initial test particle distribution using the KernelDensity Python module from the scikit-learn package (Pedregosa et al., 2011). A Gaussian kernel with a width of 0.005 was used on the semi-major axis, eccentricity and inclination; the other three angles were chosen uniformly random between 0 and 360∘. An example of this cloning in shown in Fig. 2. The black dots show the initial distribution, and the red dots are the clones. The width parameter was obtained through trial and error: I experimented with a larger width parameter, but I found that by increasing the width the eccentricity distribution of the clones, which was the most sensitive to the kernel width, became too wide.
The dynamical simulations were run using the Central Processing Unit (CPU) version of the Gravitational ENcounters with GPU Acceleration (Genga) N-body integrator (Grimm & Stadel, 2014; Grimm et al., 2022). All simulations were run on the Vega supercomputer in Slovenia using four CPU threads per simulation. All the major planets were added on their current orbits: Mercury to Neptune, plus dwarf planets Vesta and Ceres. I realise that the orbits and masses of the current planets are different from those that existed in the planet formation simulations; this procedure gravitationally shocks the orbits of the test particles. None of the terrestrial planet formation simulations, however, reproduced the terrestrial planets exactly as they are today, and I am not aware of an alternative approach for generating the initial conditions.
Simulations were run for 1 Gyr with a time step of 0.01 yr ( time steps in total). Each simulation began with 8192 test particles representing the leftover planetesimal population. Simulation data was written to disc every 100 000 years (100 kyr). Planetesimals were removed once they were further than 40 au from the Sun (whether bound or unbound), or when they collided with a planet or when they were closer than 0.05 au from the Sun. To keep simulation time reasonable I did not include corrections from General Relativity because this results in a 33% speed penalty. The code takes about 3-6 million steps per particle per second on Vega using four threads on AMD EPYC 7H12 CPUs clocked at 3.1 GHz; maximum speed was achieved when the number of test particles . I used two CPU threads when and only one when . For each terrestrial planet formation model I ran 16 simulations with a total of 131k test particles.
3 Results
3.1 Global dynamical evolution
In Fig. 3 I show snapshots of the semi-major axis versus eccentricity for the ABI simulations, where the colour coding of the planetesimals is a proxy for their inclination, and the panel titles refer to the simulation time in million years (Myr). After 400 Myr almost all Earth and Venus crossers are eliminated, but there is still a substantial population of Mars crossers. There is also a small group of planetesimals that are seemingly stable between the Earth and Mars for 1 Gyr. This is an artefact of the simulations because these bodies would have had their orbits modified by the Yarkovsky effect and therefore this region is expected to be empty – as it is today. Simulations of these same initial conditions including the Yarkovsky effect are ongoing and will be discussed in future works. Last, there is a long-term stable population at moderate inclination between Mars and the resonance, which is located at 2.05 au at low inclination (Williams & Faulkner, 1981); this population is akin to the Hungaria family (Bottke et al., 2012). Morbidelli & Nesvorný (1999) showed that bodies beyond the resonance may slowly leak onto Mars-crossing orbits due to chaotic motion induced by resonance overlap. In addition, some bodies remained on stable orbits in the main belt between 2.1 and 3 au.
A different visualisation of the evolution of the planetesimals is shown in Fig. 4, which is a heat map of the semi-major axis and perihelion distance of all bodies in all simulations for each model. The top panel is for GT, the middle for ABI, and the bottom for DD. The gradient increases from blue to red, indicated by the colourbar, with blue dots indicating few planetesimals ventured in this region of phase space, and red indicating that many remained in another region. The terrestrial planets plus Vesta and Ceres are shown with large bullets. These heat maps just show the counts per bin, and not the residence time though it is easy to convert between the two. It is immediately clear that the maps for the ABI and DD models show some similarities, while the map for the GT model is rather different.
The distributions shown in the heat maps depend on the initial conditions for the models. In the ABI and DD models, there is a sharp increase in density, or residence time, for planetesimals that are Earth-crossing versus those that are not Earth-crossing, as indicated by the two red triangular regions in these panels. In the ABI model, the first region appears at au and au, while the second one is in the DD model and has au and au. The boundaries in semi-major axis for the second region are marked by the resonance at 2 au and the 3:1 Jovian resonance at 2.5 au. In both plots the E-belt region between 1.7 au and 2 au, as proposed by Bottke et al. (2012), remains occupied for a long time. In the DD model the inner main belt is heavily populated, a consequence of the initial conditions. In contrast, the GT heat map looks substantially different, with the reddest region having au and au apart from a small sample with au and . The GT model has the greatest proportion of initial Earth-crossers, so that it is expected that this population has a shorter average lifetime than those of the ABI and DD models. The heat maps suggest that the initial perihelion distribution plays an important role in the rate of decline of the population, and thus in the impact rate onto the inner planets.
In Fig. 5 I present a snapshot of the semimajor axis-inclination distribution at the end of the simulations for all three different cases, and compare them to the actual main belt in the bottom right panel. It is clear that the DD case fares poorly in that it only produces the Hungaria group near au and and the inner main belt underneath the secular resonance for which au, but not the high inclination population above the . The GT and ABI cases populate all of the main groups to some extent, except the one between 3 au and 3.25 au. The ABI model leaves more mass in the Hungaria group than the GT case, while GT populates the middle main belt between 2.5 au and 2.8 au better than ABI does. Overall the GT and ABI models fare better than DD.
3.2 Impact chronology and population decline
Planet/Model | ||||
Earth GT | 0.13, 9.3 | 0.69, 34.2 | 0.16, 89.5 | 0.012, 303 |
Earth ABI | 0.068, 8.0 | 0.46, 36.8 | 0.46, 109 | 0.016, 351 |
Earth DD | 0.10, 6.4 | 0.43, 30.8 | 0.44, 105 | 0.021, 306 |
Mars GT | 0.12, 11.9 | 0.65, 46.4 | - | 0.23, 154 |
Mars ABI | - | 0.20, 32.2 | 0.75, 116 | 0.044, 272 |
Mars DD | 0.16, 16.8 | - | 0.64, 78.5 | 0.21, 206 |
Remain GT | 0.17, 8.5 | 0.56, 35.1 | 0.23, 88.6 | 0.016, 1470 |
Remain ABI | 0.13, 10.9 | 0.41, 42.0 | 0.45, 116 | 0.023, 1370 |
Remain DD | 0.20, 17.9 | 0.43, 65.0 | 0.27, 143 | 0.10, 5233 |
The complementary cumulative distribution of impacts of planetesimals onto the Earth and Mars, as well as the fraction of remaining planetesimals is fitted as
where the constants are weights adding up to 1, the variables are e-folding times, and e is the base of the natural logarithm; the number of terms required turned out to be usually four and sometimes three. I prefer this formulation rather than using a stretched exponential – – because the multi-exponential formulation yields a better fit, and the fitting constants yield insight about the initial conditions and fundamental dynamical processes; specifically the number of terms reveal the number of sinks or processes involved. The results of the fits are given in Table 1. For impacts on the Earth the values of and are rather similar across the models, being identical to within 25% and having values near 35 Myr and 100 Myr; the difference increase to to 37% for .
Nesvorný et al. (2017) used a similar fit with three terms to model the impact rate rather than the cumulative number of impacts. The first time constant fitted by Nesvorný et al. (2017) is 37 Myr and the second is 160 Myr. Both are of the same order as mine. Nesvorný et al. (2017) do not provide an explanation for these time constants and just state they are required to fit their high impact flux in the first Gyr. I argue that the clustering of ages around specific values for different models suggests that these time constants reflect intrinsic processes in the system, and that the fitting constants are a reflection of the initial orbital distribution of the planetesimals.
For Mars the middle two time constants in the ABI model are similar to those for the Earth, but they carry different weights; the martian impact time constants are more different from Earth’s for the other two models. For the fraction of remaining planetesimals, the same equation also works well. In particular, for the GT model the first three time constants are similar to impacts on Earth, while the weights are identical within factors of 25%. For Mars the time constants of its impact chronology for the DD model resembles the overall decline of that population. This indicates that the impact rate is modulated by the overall rate of decline, or removal, of the planetesimals from the reservoirs.
These results beg the question whether the coefficients are related to the initial fraction of Earth (and Mars) crossers. In the GT, ABI and DD models the initial fraction of Earth crossers is 70%, 50% and 35%, which are close to the coefficients for impacts on Earth. For the ABI and DD models for Earth, so the remaining particles not controlled by Earth and only by Mars impact the Earth on a timescale . For impacts on Mars only three terms are needed, and for the ABI and DD models martian impacts occur mostly on timescales and , in contrast to for the GT model. For all models the impact rate on the terrestrial planets is related to the overall decline of the whole population As such, if no planetesimals remain to impact Mars because most of them are gone after a few times , then the rate of martian impacts must also decline in tandem.
In summary, the time constants for both the impacts and the remaining fraction seem to be clustered in four groups: one group at Myr, another at Myr, a third at Myr and a fourth that is generally Myr. This clustering suggests that the rate of impacts is partially set by the rate of decline of planetesimals, which in turn is set by their leakage rate from their initial reservoirs. For comparison: Nesvorný et al. (2017) used three terms and found Myr, Myr and Myr for the rate of impacts on the Earth, even though their study only included the main asteroid belt plus E-belt subjected to an episode of giant planet migration. These similar values support my earlier hypothesis that the time constants are features of the system. Furthermore, the crater chronologies of Neukum et al. (2001) and Werner et al. (2023) have e-folding timescales of Myr, which is of the same order as , indicating that the lunar craters mostly reflect the impacts from an ancient Mars-crossing population. This result implies that the number of Earth crossers at the time that the Moon formed had to be low, probably less than half that of the Mars crossers.
3.2.1 The evolution of the perihelion distribution
In Fig. 6 I plot the evolution of the cumulative perihelion distribution for each model in the left column, and the complementary cumulative distributions of the impacts on Earth (blue) and Mars (red), as well as the fraction of planetesimals remaining (black) and the total planetesimals that are removed (orange) in the right column. The GT model is the top row, the middle is for ABI, and the bottom for DD. In all models the cumulative distribution rapidly changes shape, with the fraction of Earth crossers dropping quickly, and the fraction of Mars crossers reducing more slowly. At later stages, usually after 700 Myr, almost all of the remaining planetesimals have au. The rate of change of the distribution can be approximately quantified as follows. For all three models the fraction of Earth crossers normalised by the remaining population decreases exponentially with an e-folding time of Myr, while for the Mars crossers in the DD model the fraction decreases with an e-folding time of 220 Myr while it is increased to 430 Myr for the GT and ABI models. Thus, for the DD model the fraction of Earth crossers normalised by the remaining population declines at the same rate as that of the Mars crossers, while for the ABI and GT models the Mars crossing fraction declines much slower.
I argued that the initial perihelion distribution is the main constraint in determining the impact chronology on the Earth and Mars as well as global population decline. This argument is evident from the figure: the initial population decline and impact rate for the GT model is much more rapid than for the ABI and DD models, which is evident from the fitting constants in Table 1. For the ABI and GT models the terrestrial impacts track the global decline for the first 100 Myr and begin to deviate from there, with the normalised rate of impacts on Mars being always slower than on Earth. In contrast, the terrestrial impact chronology closely tracks all the planetesimals depletion in the GT and ABI models, while Mars’ impact chronology is initially off to a slower start but its impact rate converges after 200 Myr. Beyond this point, in all models the impact chronology is set by the dynamical evolution of the Mars crossers, which evolve on a timescale of about 80–100 Myr (Michel et al., 2000), in other words , because encounters with Mars will need to supply the planetesimals to Earth, either directly or via the .
The behaviour of the remaining fraction of planetesimals is different: it levels off after a few hundred Myr of evolution due to planetesimals being parked on long-lived orbits beyond Mars. Some of these slowly leak onto Mars-crossing orbits due to chaotic diffusion (Morbidelli & Nesvorný, 1999), and some of these planetesimals collide with the terrestrial planets.
3.3 Removal mechanism behind the fitted timescales
A greater insight into the origin of the timescales is to investigate which planetesimals are removed at what time, and from where. Figure 7 is a heat map showing the original semi-major axis and eccentricity of all planetesimals that are removed from the simulation at different time quartiles. The top left panel shows the removal for the first quartile of the population, the top right panel the second quartile, the bottom left shows the third quartile, and the bottom right shows the last quartile. The top four squares are for the GT model, and the bottom four are for the ABI model. Throughout the simulation planetesimals are removed from all over, as evidenced by the blue dots, and instead I focus on the orange regions.
For the GT model in the top left panel, which lasts for 7 Myr and is comparable to , removed planetesimals initially reside on Earth-crossing orbits with semi-major axis between 1.2 and 2.5 au. There are clear indicators of rapid removal of planetesimals beginning near 2 au and 2.5 au even at low initial eccentricity, corresponding to the and 3:1. The next quartile, shown in the top right panel, occurs between 7 Myr and 21 Myr and still has the densest part on Earth-crossing orbits but at shorter semi-major axis, between 1.1 au and 1.6 au. The high density region shifts to shorter semi-major axis in the third panel, lasting from 21 Myr to 52 Myr. Combined, these quartiles last for about 1.5. The reason for this shift towards shorter semi-major axis is that it takes longer to eliminate these objects through collision or by scattering them into the or the 3:1. As such, it appears likely that is related to the rate of removal of Earth crossers. The last panel, lasting 948 Myr, shows the slowest-declining population, which are Mars crossers with semi-major axis shorter than 2 au. This last quartile is dominated by and – although the latter has a low coefficient – and suggests that is the timescale of evolution of Mars crossers.
For the ABI model the top left panel has the highest density of planetesimals initially on Earth-crossing orbits with initial semi-major axes between 1.3 au and 2 au, with also some bodies in the – see the vertical points near 2 au at low eccentricity. This quartile lasts for 12 Myr, comparable to , suggesting that is related to removal through the and high-eccentricity Earth (and Venus) crossers. The next quartile, lasting from 12 Myr to 36 Myr, still has a large number of Earth crossers, but there are also a fair amount of Mars crossers in there; the region is still being emptied out. By now half of the planetesimals are removed. During the third quartile, lasting from 36 Myr to 91 Myr, most of the planetesimals that are removed are Mars crossers with initial semi-major axes between 1.2 au and 2 au. The rate of population decline is beginning to slow down, with the terms with Myr and Myr setting the rate, once again suggesting that is related to the removal of Earth crossers and to that of Mars crossers. The fourth quartile lasts for 900 Myr, is modulated by the terms with and at late stages, and consists mostly of Mars crossers. However, the E-belt region beyond Mars, with initial semi-major axis between 1.6 au and 2 au and perihelion au (Bottke et al., 2012), also plays a role; this region appears to play almost no role in the GT simulations.
In summary, I argue that is likely a measure of the decline of high-eccentricity Earth-crossers and bodies initially in the and 3:1, describes Earth and Venus crossers, and pertains to the Mars crossers, and possibly the E-belt. The term , which is the last few percent, is due to the decline of bodies originally beyond Mars on more stable orbits such as the inner main belt.
One way to verify the mechanisms behind and is to use the equations developed by Öpik (1951). The probability per orbital revolution of coming within a distance of the planet is given by
where is the encounter velocity scaled by the planet’s orbital velocity, and is the -component of this velocity – which is essentially the radial component of the encounter velocity. For the initial conditions considered in the simulations the value of for the Earth ranges from 1.5 to 2.2, and . The radius at which a collision occurs can be computed from energy and angular momentum conservation, and is given by ; when for the Earth and 1.13 for Venus. This leads to a probability of collision with the Earth or Venus of per orbit of a planetesimal that crosses the Earth’s or Venus’ orbit. The corresponding e-folding time for collision is then Myr using a typical semi-major axis of au. This timescale is far longer than the timescales for removal numerically computed here. In the simulations only of planetesimals are removed through collision with an inner planet; the rest are removed by ’ejection’, that is, they are being scattered into the secular resonance or 3:1 mean-motion resonance through repeated encounters with the inner planets. These resonances pump up the eccentricities of the planetesimals and they ultimately collide with the Sun or encounter Jupiter (Gladman et al., 1997).
To calculate the removal timescale I need an estimate of the cumulative change in semi-major axis due to scattering. For small deflections Arnold (1965) computes that the cumulative effect of distant encounters beyond logarithmically increases the change in energy through encounters closer than , resulting in an enhancement by a factor of 4.5. Since is large and deflections are small, I can extrapolate this relation for all encounters beyond , which increases it to 6.9. Therefore, a rough timescale of ’ejection’ of a planetesimal through repeated Earth encounters is about seven times shorter than that of removal through collision, resulting in an e-folding time of 60 Myr. By harmonically adding the effect of elimination through Venus encounters (Arnold, 1965) I arrive at an e-folding timescale of removal through combined Earth and Venus encounters of 30 Myr, which is comparable to found by my dynamical simulations.
For Mars and so that . Elimination through ’ejection’ and collisions with Mars results in an e-folding time of about 200 Myr when au, which is twice as long as found numerically here as well as in the literature (Michel et al., 2000; Fernández & Helal, 2023), but still of the same order. Objects near Mars undergo large changes in eccentricity due to the proximity of the secular resonance, which can bring them to Earth-crossing orbits which assists in their removal. Still, I argue that is associated with the removal of planetesimals by interaction with Mars.
For both models the rate of decline during the last quartile, which lasts Myr for all models, is dominated by Mars crossers, while during the second and third quartile the rate is still influenced by remaining Earth crossers. However, for the GT model Myr for the decline of the population, akin to the timescale found by Michel et al. (2000) for the Mars crossers. Given the similarity between the initial cumulative semi-major axis distribution in the ABI and DD models and the slower impact rate of the latter, it stands to reason that the initial perihelion distribution is the most important parameter controlling the impact rate. The fitting constants therefore probably represent the (initial) perihelion and semi-major axis distribution.
One final point of note pertains to the total fraction remaining after 1 Gyr. For both the GT and ABI models this is about 1%, but for the DD model this is about 10%. The last value is problematic because the current mass in the main belt is about , which probably declined by about 50%-65% over the past 4 Gyr (Minton & Malhotra, 2010; Nesvorný et al., 2017; Brasser et al., 2020), with even less than that in the inner main belt. In the DD model the remaining 10% resides in the inner main belt, implying that the initial population in planetesimals at 4.5 Ga was no greater than . This is insufficient to account for the lunar craters and its highly siderophile element abundances in its mantle (Morbidelli et al., 2018; Brasser et al., 2020). Therefore I will have to discount the initial conditions of the DD model as being viable for late accretion.
4 Spherule layer constraints on the Archean impact rate and source population
It has been suggested that terrestrial spherule beds are a record of ancient terrestrial impacts (Simonson & Glass, 2004). Johnson & Melosh (2012) used the spherule layer thickness from the recorded list of known spherule beds in Glass & Simonson (2012) to compute the diameter range of the impactor that created the spherule bed. It is not clear whether this list is complete, but I will treat it as such. In Table 2 I have listed the names and ages of the largest of these spherule beds, the average impactor diameter from the range listed in Johnson & Melosh (2012), and the corresponding number of impacts on the Earth for objects with diameter km using a size-frequency distribution slope of . For completeness I added the Sudbury layer with the proposed impactor diameter using the formula from Johnson & Melosh (2012) but I do not use it in the analysis. There is also the Vredefort crater in South Africa with an approximate age of 2 Ga, which may have left traces of a bed as far as Russia (Huber et al., 2014; Allen et al., 2022). No spherule beds have been found between 0.6 and 1.7 Ga ago, but this time interval has not been extensively explored for impact spherules (Simonson & Glass, 2004).
Bottke et al. (2012) used the presence of the spherule layers and their thickness to compute the impact rate onto the Earth from a declining population of asteroids placed just beyond the orbit of Mars. Here I will use a slightly different approach. From Table 2 I compute that there were 17k impacts on Earth between 3.47 Ga and 2 Ga with diameter km. This results in an average impact rate of 11.5 Myr-1. The uncertainties are factors of a few due to the possible range of the impactor diameters. What reservoir provided these planetesimals?
Name | Age [Ga] | [km] | |
Sudbury | 1.82 | 15 | 225 |
Grænsesø | 2.0 | 60 | 3600 |
Kuruman | 2.46 | 15 | 225 |
Dales George | 2.49 | 40 | 1600 |
Bee George | 2.54 | 14 | 196 |
Reivilio | 2.56 | 22 | 484 |
Paraburdoo | 2.57 | 22 | 484 |
Monteville | 2.62 | 38 | 1444 |
S4 | 3.24 | 43 | 1849 |
S3 | 3.25 | 56 | 3136 |
S2 | 3.26 | 48 | 2304 |
S1 | 3.47 | 41 | 1681 |
During the Archaean and later the cratering rate on the Earth can be computed analogous to the outer Solar System, and is simply (Brasser et al., 2025), where is the flux of objects on Earth-crossing orbits, and is the impact probability with the Earth. The results in Fig. 6 seem to apply that this relation is justified by the similar shape of the terrestrial and martian impact curves with those of removed planetesimals. Furthermore, at late times ( Myr of simulation time) most objects are entirely beyond Mars so that the impact rate is determined by the rate at which these objects become Mars crossers. This situation is similar to that of Kuiper Belt and Scattered Disc objects that leak into the realm of the giant planets from beyond Neptune (Duncan & Levison, 1997).
In practice the flux is equal to (Duncan et al., 1995), where is the total number of objects in all reservoirs with km. I follow Duncan et al. (1995) and compute the rate of decline of planetesimals as the number of planetesimals that left the system () divided by the final number of the particles at the end (), and the time interval (). In other words
It is worth mentioning that generally depends on the time range over which it is calculated because the rate of decline is generally not constant with time; instead, the rate of decline is a function of time itself: it decreases with increasing time. An exception is an exponential decay of the form e-t/τ, for which we have , which is constant. From the simulation output at late times . In addition, apart from the small fraction of objects that are ejected due to encounters with Jupiter, the far majority are lost due to collisions with a planet or the Sun; on the way to the Sun they are temporary Earth crossers, so that to a good approximation is equal to the rate of decline of the total remaining population (see also Fig. 6).
In the simulations, the impact probability with Earth ranges between 7.3% (ABI and DD) and 9% (GT), so I will take 7.5% as the average. This impact probability implies that 227k objects with km were lost between 3.47 Ga and 2 Ga from the various small body reservoirs in the inner Solar System. Assuming that and Gyr then the aforementioned 227k objects represent a decline of 63% by 2 Ga in the population remained at 3.47 Ga. A similar value of Gyr was also found by Nesvorný et al. (2017). If instead Myr, as suggested by the impacts on the Earth and Mars, then the population remaining at 3.47 Ga would have declined a further 99.3% by 2 Ga. Such an extreme depletion can be ruled out because it would imply an implausibly large mass in the initial population at 4.5 Ga. Using Gyr implies that there were about 617k objects with km at 3.47 Ga available to supply the terrestrial impactors. The main belt between 2.1 au and 3.3 au barely declines (Minton & Malhotra, 2010; Nesvorný et al., 2017; Brasser et al., 2020), so that these 617k planetesimals must have been in the leftover population, or in the E-belt, or both.
At 3.47 Ga the remaining population is about 1% (ABI and GT) or 9% (DD) of the population at 4.5 Ga. The remaining fraction in the DD simulations is too high, so I will use a total decline of 99% after 1 Gyr of evolution. This implies a primordial population of the order of 60 million objects with km at 4.5 Ga, which corresponds to a mass in leftover planetesimals of about 0.015 . This estimate is consistent with earlier values reported in the literature (Brasser et al., 2021; Nesvorný et al., 2022).
What about the E-belt of Bottke et al. (2012) being the source? Brasser et al. (2020) conducted dynamical simulations of the E-belt for 1 Gyr. For this work these were extended to 2 Gyr. For the E-belt the impact probability with Earth is , requiring 340k bodies to leak from the E-belt between 3.47 Ga and 2 Ga to supply all of the terrestrial impacts. The fraction remaining can be fit by , so that at late times Myr-1. In the simulations the E-belt loses an additional 93% of its mass between 3.47 Ga and 2 Ga, so that I can safely assume that at 3.47 Ga the E-belt had 340k objects with km. The simulation output shows that at 3.47 Ga the E-belt only has 3.9% of its mass at 4.5 Ga, implying that the primordial number of E-belt objects with km is approximately 8.7M assuming no collisional evolution. The required mass for this many objects is about 0.002 , which is a factor of a few higher than what was suggested by Bottke et al. (2012). It is therefore more likely that the Archean impacts that created the terrestrial spherule beds were predominantly () supplied by leftover planetesimals on meta-stable orbits slowly leaking onto Earth-crossing orbits from beyond Mars.
5 Conclusions
I have simulated the dynamical evolution of leftover planetesimals for 1 billion years from three different models of terrestrial planet formation: the Grand Tack (Walsh et al., 2011), the Depleted Disc (Izidoro et al., 2014), and implantation (Brasser, 2025). Each one of these models yields a different initial distribution of orbital elements for the leftover planetesimals. In particular, the initial fraction of Earth crossers for the GT is twice as high as for DD, which has implications for the dynamical evolution and impact rate onto the Earth and Mars. The DD model leaves too much mass after 1 Gyr to be consistent with the current mass of the main asteroid belt.
I fit the cumulative number of impacts and the remaining fraction of planetesimals with a function that is a sum of three to four exponentials whose e-folding timescales cluster around 10 Myr, 35 Myr, 100 Myr and 200 Myr. These distinct timescales are the result of different processes removing the planetesimals at different rates, with high-eccentricity Earth crossers and planetesimals in the secular resonance being removed first (), followed by Earth crossers (), Mars crossers () and finally objects leaking on to Mars crossers from the E-belt and the inner main belt (). The timescale is of the same order as that of the lunar chronology of Neukum et al. (2001) and Werner et al. (2023), suggesting that the lunar craters are sourced predominantly from an ancient population of Mars crossers, and that the number of Earth crossers at the time that the Moon formed had to be rather low. As such, the initial perihelion distribution of the leftover planetesimals plays a decisive role in the overall rate of decline of the population and the impact rates onto the Earth and Mars. These results place constraints on models of terrestrial planet formation, wherein the number of Earth crossers at the time of the Moon’s formation had to be relatively low compared to the Mars crossers.
To calibrate the impact rate at late times I use the Archean spherule beds as listed in Johnson & Melosh (2012), converting the average diameter of the impactor to the number of objects with diameter km striking the Earth over 1.47 Gyr between 3.47 Ga and 2 Ga. I conclude that the most likely source of these impacts are the leftover planetesimals rather than the main belt or the E-belt, and that this Archean flux requires a primordial leftover population with a mass of about 0.015 at 4.5 Ga. Implications for the lunar cratering rate and the amount of mass accreted by the Moon and Mars will be presented in the future.
Acknowledgements.
I thank Stephanie Werner and Emily W. Wong for comments. I acknowledge EuroHPC Joint Undertaking for awarding the project EHPC-REG-2024R01-093 access to Vega at IZUM, Slovenia. This study is supported by the Research Council of Norway through its Centres of Excellence funding scheme, project No. 332523 PHAB. GENGA can be obtained from https://bitbucket.org/sigrimm/genga/.References
- Allen et al. (2022) Allen, N. H., Nakajima, M., Wünnemann, K., Helhoski, S., & Trail, D. 2022, Journal of Geophysical Research: Planets, 127, e2022JE007186
- Arnold (1965) Arnold, J. R. 1965, ApJ, 141, 1536
- Bottke et al. (2007) Bottke, W. F., Levison, H. F., Nesvorný, D., & Dones, L. 2007, Icarus, 190, 203
- Bottke et al. (2002) Bottke, W. F., Morbidelli, A., Jedicke, R., et al. 2002, Icarus, 156, 399
- Bottke et al. (2012) Bottke, W. F., Vokrouhlický, D., Minton, D., et al. 2012, Nature, 485, 78
- Brasser (2025) Brasser, R. 2025, A&A, 694, A318
- Brasser et al. (2016) Brasser, R., Matsumura, S., Ida, S., Mojzsis, S. J., & Werner, S. C. 2016, ApJ, 821, 75
- Brasser et al. (2021) Brasser, R., Mojzsis, S. J., Werner, S. C., & Abramov, O. 2021, Icarus, 361, 114389
- Brasser et al. (2020) Brasser, R., Werner, S. C., & Mojzsis, S. J. 2020, Icarus, 338, 113514
- Brasser et al. (2025) Brasser, R., Wong, E. W., & Werner, S. C. 2025, A&A, 695, A276
- Chambers (2001) Chambers, J. E. 2001, Icarus, 152, 205
- Day et al. (2012) Day, J. M. D., Walker, R. J., Qin, L., & Rumble, III, D. 2012, Nature Geoscience, 5, 614
- Duncan & Levison (1997) Duncan, M. J. & Levison, H. F. 1997, Science, 276, 1670
- Duncan et al. (1995) Duncan, M. J., Levison, H. F., & Budd, S. M. 1995, AJ, 110, 3073
- Farinella & Vokrouhlický (1999) Farinella, P. & Vokrouhlický, D. 1999, Science, 283, 1507
- Farinella et al. (1998) Farinella, P., Vokrouhlický, D., & Hartmann, W. K. 1998, Icarus, 132, 378
- Fernández & Helal (2023) Fernández, J. A. & Helal, M. 2023, Icarus, 394, 115398
- Gladman et al. (1997) Gladman, B. J., Migliorini, F., Morbidelli, A., et al. 1997, Science, 277, 197
- Glass & Simonson (2012) Glass, B. P. & Simonson, B. M. 2012, Elements, 8, 43
- Gomes et al. (2005) Gomes, R., Levison, H. F., Tsiganis, K., & Morbidelli, A. 2005, Nature, 435, 466
- Granvik et al. (2018) Granvik, M., Morbidelli, A., Jedicke, R., et al. 2018, Icarus, 312, 181
- Grimm & Stadel (2014) Grimm, S. L. & Stadel, J. G. 2014, ApJ, 796, 23
- Grimm et al. (2022) Grimm, S. L., Stadel, J. G., Brasser, R., Meier, M. M. M., & Mordasini, C. 2022, ApJ, 932, 124
- Hansen (2009) Hansen, B. M. S. 2009, ApJ, 703, 1131
- Huber et al. (2014) Huber, M. S., Črne, A. E., McDonald, I., et al. 2014, Geology, 42, 375
- Izidoro et al. (2014) Izidoro, A., Haghighipour, N., Winter, O. C., & Tsuchida, M. 2014, ApJ, 782, 31
- Johnson & Melosh (2012) Johnson, B. C. & Melosh, H. J. 2012, Nature, 485, 75
- Lammer et al. (2021) Lammer, H., Brasser, R., Johansen, A., Scherf, M., & Leitzinger, M. 2021, Space Sci. Rev., 217, 7
- Mah & Brasser (2021) Mah, J. & Brasser, R. 2021, Icarus, 354, 114052
- Michel et al. (2000) Michel, P., Migliorini, F., Morbidelli, A., & Zappalà, V. 2000, Icarus, 145, 332
- Migliorini et al. (1998) Migliorini, F., Michel, P., Morbidelli, A., Nesvorny, D., & Zappala, V. 1998, Science, 281, 2022
- Minton & Malhotra (2010) Minton, D. A. & Malhotra, R. 2010, Icarus, 207, 744
[…]
- Morbidelli & Nesvorný (1999) Morbidelli, A. & Nesvorný, D. 1999, Icarus, 139, 295
- Morbidelli et al. (2018) Morbidelli, A., Nesvorný, D., Laurenz, V., et al. 2018, Icarus, 305, 262
- Morbidelli et al. (2012) Morbidelli, A., Tsiganis, K., Batygin, K., Crida, A., & Gomes, R. 2012, Icarus, 219, 737
- Nesvorný & Morbidelli (1998) Nesvorný, D. & Morbidelli, A. 1998, AJ, 116, 3029
- Nesvorný et al. (2017) Nesvorný, D., Roig, F., & Bottke, W. F. 2017, AJ, 153, 103
- Nesvorný et al. (2022) Nesvorný, D., Roig, F. V., Vokrouhlický, D., et al. 2022, ApJ, 941, L9
- Neukum et al. (2001) Neukum, G., Ivanov, B. A., & Hartmann, W. K. 2001, Space Sci. Rev., 96, 55
- O’Brien et al. (2006) O’Brien, D. P., Morbidelli, A., & Levison, H. F. 2006, Icarus, 184, 39
- Öpik (1951) Öpik, E. J. 1951, Pattern Recognition and Image Analysis, 54, 165
- Pedregosa et al. (2011) Pedregosa, F., Varoquaux, G., Gramfort, A., et al. 2011, Journal of Machine Learning Research, 12, 2825
- Raymond et al. (2009) Raymond, S. N., O’Brien, D. P., Morbidelli, A., & Kaib, N. A. 2009, Icarus, 203, 644
- Raymond et al. (2006) Raymond, S. N., Quinn, T., & Lunine, J. I. 2006, Icarus, 183, 265
- Simonson & Glass (2004) Simonson, B. M. & Glass, B. P. 2004, Annu. Rev. Earth Planet. Sci., 32, 329
- Vokrouhlický & Farinella (1998) Vokrouhlický, D. & Farinella, P. 1998, AJ, 116, 2032
- Vokrouhlický & Farinella (1999) Vokrouhlický, D. & Farinella, P. 1999, AJ, 118, 3049
- Vokrouhlický & Farinella (2000) Vokrouhlický, D. & Farinella, P. 2000, Nature, 407, 606
- Walsh & Levison (2019) Walsh, K. J. & Levison, H. F. 2019, Icarus, 329, 88
- Walsh et al. (2011) Walsh, K. J., Morbidelli, A., Raymond, S. N., O’Brien, D. P., & Mandell, A. M. 2011, Nature, 475, 206
[…]
- Werner et al. (2014) Werner, S. C., Ody, A., & Poulet, F. 2014, Science, 343, 1343
- Wetherill (1977) Wetherill, G. W. 1977, Lunar and Planetary Science Conference Proceedings, 1, 1
- Williams & Faulkner (1981) Williams, J. G. & Faulkner, J. 1981, Icarus, 46, 390
- Woo et al. (2021) Woo, J. M. Y., Grimm, S., Brasser, R., & Stadel, J. 2021, Icarus, 359, 114305